An oil drop of mass 10-12 gm is floating on the free surface of a liquid. At any instant the position of the drop can be determined within the error of 10-4 cm. what will be the error in measurement of its velocity.
1
Expert's answer
2019-11-21T10:07:04-0500

The speed is defined as the ratio of distance to time, then the error will be determined as follows:

"\Delta v= \frac{t\cdot \Delta s + s\cdot \Delta t}{t^{2}}"

Since there are no exact speed values, it is impossible to accurately determine the error.